My question is what key words should I look for (and translating the thing):
Yissania and Charlotte walked on treadmills at the same time at the gym. Yissania walks at 5 mph while Charlotte walks 4.2 mph. If Yissania walked .6 miles farther than charlotte, how long did they use the treadmills.
I appreciate the help,
John

Found 2 solutions by stanbon, noxie:
Answer by stanbon(75887)

Yissania and Charlotte walked on treadmills at the same time at the gym. Yissania walks at 5 mph while Charlotte walks 4.2 mph. If Yissania walked .6 miles farther than charlotte, how long did they use the treadmills.
-----
Yissania DATA:
rate = 5 mph : distance = x+0.6 miles ; time = d/r = (x+0.6)/5 hrs
-----------------
Charlotte DATA:
rate = 4.2 mph ; distance = x miles ; time = d/r = x/4.2 hrs
-----
Equation:
time = time
(x+0.6)/5 = x/4.2
----------------------
4.2x + 4.2*0.6 = 5x
0.8x = 2.52
x = 3.15 hrs (time each used the treadmill)
3.15 hrs = 3 hrs 9 minutes
==================================
Cheers,
Stan H.
=====================

Answer by noxie(1)

The answer above is not correct. 3.15 is the distance that Charlotte has traveled. To get the time they both traveled, you would take the 3.15 and divide it by Charlotte's rate, which is 4.2 mph
3.15/4.2= time traveled
3.15/4.2 = .75 hours. .75 hours is the final and correct answer to this question. (I know this because the textbook has 3/4 of an hour listed as the answer and I got it doing this. The person above confused themselves and wrote hours instead of mph)